OpenStudy (anonymous):

8 ft 6 in - 3 ft 11 in

OpenStudy (anonymous):

i have possible answers if you need them

OpenStudy (anonymous):

again 1 ft = 12 inches now do the math 8.6 3.11 ---- 4.5

OpenStudy (anonymous):

1ft=12 in 8 ft = 12*8=96 in 96+6 =102 in and 3ft 11 in = 47 in so 102-47= 55 in 55 in = 4 ft 7 in

OpenStudy (anonymous):

love men play hard you were right

OpenStudy (anonymous):

awh thanks ;)